OpenStudy (anonymous):

(How Much 25% Acid Solution Should Be Mixed With 85% Solution In Order To Obtain 300 Liters Of 65% Solution)

OpenStudy (amistre64):

a + b = t ; *-x xa + yb = zt -xa -xb = -xt xa +yb = zt ------------- b(y-x) = t(z-x) b@y = t(z-x)/(y-x) amount@25% = 300(.65-.85)/(.25-.85)